OpenStudy (anonymous):

whats 46/17 as inproper fraction (in lowest terms

OpenStudy (solomonzelman):

` F O R M U L A ─ ` \(\LARGE\color{blue}{ \rm A\frac{b}{c}=\frac{(A \times c)+b}{c}}\) `Hint: ` \(\large\color{blue}{ \rm46= (17\times 2)+12 }\)